Please show me how to figure this out...
I need to find the x-value of
x-2y=-6
y=3x-2
A)x=1
B)x=2
C)x=4
D)x=4/5
thanks!

Answer by vidhyak(80)

x-2y=-6
y=3x-2

Lets rewrite the equations so that all variables are on one side and constant on the other side
x - 2y = -6 ----- equation 1
3x - y = 2 ----- equation 2

Multiply eq 2 by2 and subtract from eq 1
x - 2y = -6
-(6x - 2y = 4)

x - 2y = -6
-6x + 2y = -4

Solving we get -5x = -10, so x =2
Answer is B